OpenStudy (anonymous):

An object dropped from a height of 150 meters hits the ground with velocity_______m/s

OpenStudy (anonymous):

what s the height

OpenStudy (anonymous):

150 meters

OpenStudy (cwrw238):

use one of the equations of constant acceleration here a = 9.81 ms^-2 ( acceleration due to gravity), s = 150, u = 0 and v = ? use v^2 = u^2 + 2as

OpenStudy (anonymous):

So with that I got 3.1237. Would that be the answer or would it be the sqrt(3.1237)?

OpenStudy (cwrw238):

v^2 = 0 + 2 * 9.81 * 150

OpenStudy (cwrw238):

= 2943 v = sqrt 2943 m/s

OpenStudy (anonymous):

Okay thank you. I have gotten that answer before but the online program we are using says that it is wrong, so I wasn't sure I was doing it right. Thanks for the help! I am going to email the instructor to see what the problem is. Thanks again!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

After talking with the instructor, the only thing wrong it that it was actually -sqrt(2943) because the object was falling.